The following conversion from y-intercept to standard form is incorrect. Explain where the error(s) occurs and how to correct the mistake(s).
Y=1/4x+5
4(y)=4(1/4)x+5
4y=1x+5
4y-x=x-x+5
4y-x=5
4y-x-5=5-5
-x+4y-5=0

Found 2 solutions by josgarithmetic, MathTherapy:
Answer by josgarithmetic(39617)
Answer by MathTherapy(10552)


The following conversion from y-intercept to standard form is incorrect. Explain where the error(s) occurs and how to correct the mistake(s).
Y=1/4x+5
4(y)=4(1/4)x+5
4y=1x+5
4y-x=x-x+5
4y-x=5
4y-x-5=5-5
-x+4y-5=0
The error is on the 2nd LINE. 
Now, YOU need to find the error and correct it!!